# TVBench: Redesigning Video-Language Evaluation

### v1

@@ A.3 TVBench qualitative examples
This section provides qualitative examples for the different TVBench tasks. As shown in Fig.7, 8, and 9, temporal information from the input video is indispensable for correctly answering the given questions.
Figure 7: TVBench: Samples from our benchmark (1). Row 1-2: Action Antonym; Row 3-4: Action Count; Row 5-6: Action Localization; Row 7-8: Action Sequence.
Figure 8: TVBench: Samples from our benchmark (2). Row 1-2: Object Count; Row 3-4: Moving Direction; Row 5-6: Object Shuffle.
Figure 9: TVBench: Samples from our benchmark (3). Row 1-2: Scene Transition; Row 3-4: Unexpected Action.
Appendix B Spatial/Textual bias in MVBench
@@
Fig. 10-21 contain several evaluation examples from MVBench that suffer from the issues analyzed in Sec. 3. This qualitative analysis complements the quantitative results discussed in Sec 6 showing that these issues affect a large portion of MVBench.

### v2

@@ A.4 TVBench qualitative examples
This section provides qualitative examples for the different TVBench tasks. As shown in Fig.7-13, temporal information from the input video is indispensable for correctly answering the given questions.
A.5 Spatial and Textual Bias in MVBench
@@
Fig. 14-25 contain several evaluation examples from MVBench that suffer from the issues analyzed in Sec. 3. This qualitative analysis complements the quantitative results discussed in Sec 6 showing that these issues affect a large portion of MVBench.
Figure 7: TVBench: Samples from our benchmark (1). Row 1-5: Action Antonym; Row 6-7: Action Count.
Figure 8: TVBench: Samples from our benchmark (2). Row 1-3: Action Count; Row 4-6: Action Localization.
Figure 9: TVBench: Samples from our benchmark (3). Row 1: Action Localization; Row 2-6: Action Sequence; Row 7: Object Count.
Figure 10: TVBench: Samples from our benchmark (4). Row 1-4: Object Count; Row 5-6: Moving Direction.
Figure 11: TVBench: Samples from our benchmark (5). Row 1-3: Moving Direction; Row 4-6: Object Shuffle.
Figure 12: TVBench: Samples from our benchmark (6). Row 1-2: Object Shuffle; Row 3-6: Scene Transition.
Figure 13: TVBench: Samples from our benchmark (7). Row 1: Scene Transition; Row 2-6: Unexpected Action.